OpenStudy (anonymous):

i wish to make 35 using exactly 4 4's we can use any mathematical operator

OpenStudy (anonymous):

for example to make 1 = 4-4 + (4/4)

OpenStudy (anonymous):

impossible :-P

OpenStudy (turingtest):

I concur!

OpenStudy (anonymous):

i know till 34 but got real stuck at 35..

OpenStudy (anonymous):

31 was also a real challenge if u want to give it a try

OpenStudy (anonymous):

all odd nos. till 34 could be represented with 4 4's then why not 35?

OpenStudy (anonymous):

25 = 4! + \[\sqrt{4}\] - 4/4

OpenStudy (mr.math):

\(4!+\frac{44}{4}\).

OpenStudy (anonymous):

\[34 = 4! + 4 + 4 + \sqrt(4)\] \[35 = 4! + \frac {4! - \sqrt{4}}{\sqrt{4}}\]

OpenStudy (anonymous):

woahh... super mr math never thought of 44..

OpenStudy (turingtest):

wow! is that allowed?

OpenStudy (turingtest):

I specifically ignored that b/c I thought that was against the rules, but I never would have though of factorials

OpenStudy (turingtest):

thought*

OpenStudy (mr.math):

@TuringT: Your medals always level me up! :)

OpenStudy (turingtest):

They're special like that :)

OpenStudy (anonymous):

If we use 44, can we also use 0.4?

OpenStudy (anonymous):

0.4 has a decimal .. 44 uses only 4's . but we can always bend some rules ;-D

OpenStudy (anonymous):

0.4 has a decimal but sqrt(4) has a sqrt and 4! has a factorial just some food for thought ...

OpenStudy (mr.math):

mever has got a point!

OpenStudy (turingtest):

actually I think sqrt is against the rules since it is\[x^{1/2}\]

OpenStudy (mr.math):

4!=4*3*2 :P

OpenStudy (turingtest):

Well this is less defined than i thought :P

OpenStudy (anonymous):

oh .. i dont kow what to say..

OpenStudy (anonymous):

we have to use math operators and 4's

OpenStudy (anonymous):

I think if we do not allow square roots and factorials, then the puzzle becomes very limited ... I think you can reach 100 with these rules. If you allow 0.4, then you can reach higher. Since it is just a game, set the rules with your friends and stick to them while having fun ...

OpenStudy (anonymous):

! and sqrt are math operators ... Is the decimal a math operator ...?

OpenStudy (turingtest):

if math operators are just */+ and - then it can't be don I guess

OpenStudy (turingtest):

oh, and ^

OpenStudy (anonymous):

is decimal a math operator??

OpenStudy (turingtest):

no way

OpenStudy (turingtest):

operators relate numbers together, right? decimal points just specify the number.

OpenStudy (anonymous):

More food for thought: an operation is an action or procedure which produces a new value from one or more input values